Margaret bought 150 kg of beans of which 20% were red beans, 10% were green beans and the rest were black beans. She changed her order such that the percentage of red beans increased to 30% and the order of black beans was reduced by 20%. What percentage of green beans did she have after she changed the order...?..
Found 2 solutions by Theo, josgarithmetic:
Answer by Theo(13342)

total amount of beans was 150 kilograms.
20% were red.
10% were green.
the rest were black.
that gets you
.20 * 150 = 30 kilograms of red
.10 * 150 = 15 kilograms of green
150 - 145 = 105 kilograms of blackshe
she changed her order such that the percentage of red beans increased to 30% (not by 30%) and the order of black beans was reduced by 20% (not to 20%).
based on the above statement, the following occurred:
the amount of red beans became .30 * `150 = 45 kilograms.
the amount of black beans became 105 - .20 * 105 = 84 kilograms.
the amount of green beans became 150 minus 45 minus 84 = 21 kilograms.
the percent of green beans became 21 / 150 = .14 = 14%.
your solution is:
the percentage of green beans became 14% after the change in the order.




Answer by josgarithmetic(39631)

New percent of the green beans, 14%.

        P  E  R  C  E  N  T  S
        ORIGINAL   CHANGED-TO

RED      20         30

GREEN    10        100-56-30=14

BLACK    70         70-(1/5)70=56

totals  100%        100%